OpenStudy (anonymous):

Which of the following is true of an isothermal process? A. Temperature is constant. B. ΔU = 0 C. Volume is constant. D. W = Q Can be multiple answers

OpenStudy (priyar):

1."isothermal" means same temperature..so temperature doesn't change. 2.internal energy changes with temperature..so if temp. doesn't change, internal energy will also not change. 3.also use first law of thermodynamics to decide upon the last option.. so what will be the answers.. ? @Schoolninja7

OpenStudy (anonymous):

A and D

OpenStudy (priyar):

how?

OpenStudy (priyar):

how D?

OpenStudy (anonymous):

It's not b or d I meant a and c

OpenStudy (priyar):

no..thats wrong... how did u arrive at this answer...read the abv text carefully

OpenStudy (priyar):

@Schoolninja7 ?

OpenStudy (anonymous):

I looked up that ∆U=W is isothermic so b is wrong I t guessed for b an d so if their wrong a is the only answer?

OpenStudy (priyar):

"∆U=W is isothermic " this is wrong...

OpenStudy (priyar):

temp is const. which means ∆U is zero. first law of thermodynamics is ∆U=W +q so 0=W+q so -W=q or W=q (since sign convention differs).

OpenStudy (priyar):

therefore all are correct except C. any doubts?

OpenStudy (frostbite):

Perfect explanation by @priyar, just remember the internal energy of a real gas is also a function of volume, so we here assume we work with an ideal gas.